OpenStudy (babynini):

The 20th term of an arithmetic sequence is 101, and the common difference is 3. Find a formula for the nth term.

OpenStudy (babynini):

@ganeshie8 are you free? :)

OpenStudy (babynini):

hrm, from the info given I have a_20=a+(20-1)d a_20=a+(19)*3 a_20=a+57 yeah?

OpenStudy (babynini):

@jim_thompson5910 :)

OpenStudy (babynini):

we know a_20 so it comes to 101=a+19(3) a=44 so the formula for the nth term will be a_n=44+(n-1)(3) ?

OpenStudy (jtvatsim):

You've put a lot of work into this already, so let's test your equation to verify it.

OpenStudy (jtvatsim):

Testing a_20 seems like a good idea. Do we get the right answer?

OpenStudy (jtvatsim):

And.... we do!

OpenStudy (jtvatsim):

Looks like you've got it. Congrats!

OpenStudy (babynini):

yay! :)

OpenStudy (babynini):

thanks for checking ^-^ (how did you test a_20?)

OpenStudy (jtvatsim):

Just plugged in n = 20.

OpenStudy (jtvatsim):

I hoped it would be 101 as the answer, and it was.

OpenStudy (babynini):

oh haha I see. Fabulous.

OpenStudy (jtvatsim):

Of course, I needed to see that n = 1 gave me 44 as claimed, but that was kind of obvious since that's what you built the formula to do. :)

OpenStudy (babynini):

Indeed :P